# Generalised height pairings and the Albanese kernel

### v1

@@ Abstract.
The Chabauty–Coleman–Kim method in depth two describes the rational points on a curve in terms of a generalisation of Nekovář’s $p$-adic height pairing which replaces $\mathbb{G}_{m}$ with a higher Chow group. It is unclear both what the domain of definition of this pairing is, and how to compute it. This paper explores the relavence of the Beilinson–Bloch conjectures to this problem. In particular, it is shown that if $X$ is a smooth projective curve and the Albanese kernel of $X\times X$ is torsion, then there is an algorithm to compute the generalised height pairing on a pair of rational points on the Jacobian. This leads to the consideration of certain ‘motivic refinements’ of the nonabelian cohomology varieties which arise in nonabelian Chabauty.
@@ 1. Introduction
In Kim’s nonabelian generalisation of the Chabauty–Coleman method, one is instead interested in iterated extensions (henceforth referred to as mixed extensions) of Galois representations which come from geometry. The graded pieces of these iterated extensions are quotients of tensor powers of the Tate module of the Albanese of the original variety $X$. The utility of working with this space of iterated extensions is that, while the space of geometric extensions of $\mathbb{Q}_{p}$ by $H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(1))$ may be large, conjectures on the dimensions of global Galois cohomology groups predict the space of geometric extensions of $\mathbb{Q}_{p}$ by higher tensor powers are ‘small’, and in particular most local iterated extensions do not come from global iterated extensions, giving an obstruction to a $p$-adic point coming from a rational points.
More precisely, let $b\in X(\mathbb{Q})$ and let $U_{n}$ be the maximal $n$-unipotent quotient of the $\mathbb{Q}_{p}$-unipotent completion of the étale fundamental group of $X_{\overline{\mathbb{Q}}}$ with basepoint $b$. Then one defines $X(\mathbb{Q}_{p})_{n}\subset X(\mathbb{Q}_{p})$ to be the set of points extending to an adelic point $(x_{v})\in X(\mathbb{A})$ whose class in $\prod_{v}H^{1}(\mathbb{Q}_{v},U_{n})$ lies in the image of $H^{1}(\mathbb{Q},U_{n})$.
@@ 1.1. Applications to quadratic Chabauty
In full-fat quadratic Chabauty, one assigns to each pair of rational points on a smooth projective variety $X$ a Galois representation which is filtered with graded pieces $\mathbb{Q}_{p},V:=H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(1))$ and $\mathrm{Coker}(H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p})^{\otimes 2}\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}H^{2}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}))^{*}$. We write $M_{g}(\mathbb{Q};\mathbb{Q}_{p},V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$ for the set of such iterated extensions (see §2.1). This is a nonabelian cohomology variety which is slightly larger than the Selmer scheme considered by Kim. However, working with this larger space has benefits: it behaves more like the Selmer scheme of a elliptic curve minus the origin with tangential basepoint studied by Kim in [Kim10]. The reciprocity law giving a condition for a collection of local mixed extensions to come from a global mixed extension is then the generalised height pairing introduced in [BD21]. This has been used in [BD21] to determine rational points on bihyperelliptic curves of the form $y^{2}=x^{6}+ax^{4}+ax^{2}+1$, and in [Dog23] to determine the rational points on the hyperelliptic curve $y^{2}-y=x^{5}-x$.
@@
To explain why this is a difficult problem, consider the cases of classical Chabauty–Coleman and skimmed quadratic Chabauty. In the case of classical Chabauty, one determines the Chabauty–Coleman locus $X(\mathbb{Q}_{p})_{1}$ by finding a set of zero-cycles on $X$ generating a finite index subgroup of $J(\mathbb{Q})$. In skimmed quadratic Chabauty, one needs to determine the values of the $p$-adic height pairing. By analogy with the abelian case, one might hope that one can do this by evaluating the $p$-adic height function on points over number fields. However it is far from obvious that this works: for example in the case of integral points on a rank $=$ genus hyperelliptic curve, one can determine the $p$-adic height pairing from $K$-points on $X$ (for $K/\mathbb{Q}$ a Galois extension, possibly infinite) if and only if $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$ is contained in the subgroup of $(\mathrm{Sym}^{2}J(K)\otimes\mathbb{Q})^{\operatorname{Gal}(K|\mathbb{Q})}$ generated by $(x-\infty)^{2}$, for $x\in X(K)$ and $\infty$ a Weierstrass point.
In practice one instead computed the $p$-adic height pairing using a more general construction – e.g. using algorithms for computing local height pairings between divisors $D_{1}$ and $D_{2}$ with disjoint support. From the perspective of mixed extensions, this can be interpreted as constructing all mixed extensions with graded pieces $\mathbb{Q}_{p},V,\mathbb{Q}_{p}(1)$ from subquotients of Galois representations of the form $H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}}-|D_{1}|;|D_{2}|)(1)$ (see [Nek93], [Sch94]).
@@ 1.2. The Albanese kernel
This is of a piece with one of the original visions of the Chabauty–Coleman–Kim method (or “recurring fantasies”, in the language of the Beilinson–Bloch conjectures) that the Bloch–Kato conjectures should imply the existence of an algorithm for determination of the rational points on curves [Kim12]. As is pointed out in loc. cit., by analogy with the Birch–Swinnerton-Dyer conjecture one would actually hope for an algorithm whose termination is conditional on these conjectures, meaning that one nevers needs to verify the Beilinson–Bloch conjectures for a variety in order to run the algorithm, but if the conjectures are false then the process is not guaranteed to terminate.
The present paper can be seen as an elaboration on one special case of that idea. [^2] Recall that Chabauty methods are essentially never known to give algorithms for determining rational points, because of well documented problems with possible nonrational points in $X(\mathbb{Q}_{p})_{n}$ or multiplicities of zeroes of the relavent $p$-adic power series (see [BS10] for a comprehensive description). Namely, we show that the problem of determining equations for $X(\mathbb{Q}_{p})_{2}$ essentially reduces to verifying (or more accurately, exhibiting) that certain elements of the Albanese kernel of $X\times X$ are torsion. To explain this, let $h$ denote the generalised height on $X$. This may be viewed either as a function on $X(\mathbb{Q})$, or as a function on $H^{1}_{f}(\mathbb{Q},V)\times H^{1}_{f}(\mathbb{Q},V^{*}\otimes(\wedge^{2}V/\mathbb{Q}_{p}(1)))$, or as a function on $J(\mathbb{Q})\times\mathrm{CH}^{2}(X^{3})_{0}$. As explained below, the relation between these is via the cycle class map
$\mathrm{Jac}(X)(K)\times\mathrm{CH}^{2}(X^{3})_{0}\to H^{1}(K,V)\times\to H^{1}(K,V^{*}\otimes\wedge^{2}V/\mathbb{Q}_{p}(1))$
@@ 1.3. Motivic avatars
(1) $1\to\mathbb{G}_{m}\to P^{\times}\to J\times J\to 1$
@@
One can extend the Poincaré torsor to describe semi-skimmed quadratic Chabauty. One can think of this as involving working with $T$-torsors $\widetilde{P}^{\times}$ on $J$, or other abelian varieties, where $T$ is a nonsplit torus, or alternatively as simply working with the Poincaré torsor over number fields. However, beyond semi-skimmed quadratic Chabauty one would not expect any kind of motivic avatar of the Selmer scheme. A precise statement in this regard is the recent paper of Rogov which proves that when $n\geq 3$, Hain’s depth $n$ higher Albanese manifold (for a hyperbolic curve $X/\mathbb{C}$) is not algebraizable [Rog25]. An archetypal example of the kind of non-algebraic structure that can arise is that of ‘mixed Tate quadratic Chabauty’ – i.e. the study of $X(\mathbb{Z}_{p})_{2}$ when $X$ is $\mathbb{P}^{1}-\{0,1,\infty\}$. Here the relavent bilinear form comes from the $p$-adic dilogarithm function. Then there is a precise analogue of the exact sequence (1) defined in terms of the abelian category of mixed Tate motives. However, from a computational perspective, arguably a more useful analogue is the related (but slightly different) Bloch–Suslin complex.333perhaps a strict analogue would be a short exact of pointed sets constructed from the Bloch–Suslin complex, see 5.2.
In classical Chabauty–Coleman, the undetermined coefficients defining $X(\mathbb{Q}_{p})_{1}$ can be solved by evaluating the $p$-adic logarithm on a Mordell–Weil basis (or any finite index subgroup of $X(\mathbb{Q}_{p})_{1}$. In skimmed quadratic Chabauty, the undetermined coefficients defining $X(\mathbb{Q}_{p})_{U_{2}^{T}}$ can be solved by computing the cyclotomic $p$-adic height pairing on a Mordell–Weil basis (together with several other computations, e.g. local constants at primes of bad reduction, $p$-adic logarithms of diagonal cycles, etc). In semi-skimmed quadratic Chabauty, the computation is essentially the same, except that one considers all $p$-adic height pairings (i.e. for all $p$-adic idele class characters). When one considers mixed Tate quadratic Chabauty, an interesting new phenomenon arises: one needs to exhibit that certain elements of $K_{2}^{M}(\mathbb{Q})$ is torsion. In full-fat quadratic Chabauty the same phenomenon arises, with $K_{2}^{M}(\mathbb{Q})$ replaced by the Albanese kernel of $X\times X$.
@@
full-fat quadratic Chabauty $\left(\begin{array}[]{ccc}1&0&0\\ &\rho_{V}&0\\ &*&\rho_{V}^{\otimes 2}\end{array}\right)$ motivic higher Albanese set $T(X^{2})$
@@
To elaborate briefly on the second point: Dan–Cohen, Wewers, Corwin [DCW15] [DCC18] and others have used the Tannakian category of mixed Tate motives to great effect in the study of Chabauty–Kim for the $S$-unit equation. However, whilst the existence of such a category greatly clarifies structural properties, as far as the author is aware one still needs a geometric realisation of these mixed Tate motives in order to compute their $p$-adic periods. However, perhaps using an explicit description of the Galois group of mixed Tate motives in terms of algebraic cycles would give a practical approach to this computational problem [Blo91].
@@ Acknowledgements
It is a pleasure to be able to thank Minhyong Kim for the opportunity to spend so much time in the mathematical world he has created, and for his generosity and optimism. This paper benefited from many discussions with Carl Wang-Erickson about mixed extensions.
@@ 2.1. Mixed extensions
Given objects $V_{0},\ldots,V_{n}$ in an abelian category $\mathcal{C}$, we can define a mixed extension with graded pieces $V_{0},\ldots,V_{n}$ to be an object $W$ with a filtration
$W=W_{0}\supset W_{1}\supset\ldots\supset W_{n+1}=0$
together with isomorphisms $\psi_{i}:V_{i}\simeq W_{i}/W_{i+1}$. We define a morphism of mixed extensions $(W,(W_{i}),(\psi_{i}))\to(W^{\prime},(W^{\prime}_{i}),(\psi^{\prime}_{i}))$ to be a morphism $W\to W^{\prime}$ in $\mathcal{C}$, respecting the filtrations and commuting with the $\psi_{i}$ and $\psi_{i}^{\prime}$. If $\mathcal{C}$ is the category of continuous $\mathbb{Q}_{p}$-representations of a group $G$, then we can a group $U(V_{0},\ldots,V_{n})$ of unipotent automorphisms444here we view $\oplus_{i=0}^{n}V_{i}$ as a mixed extension of the $V_{i}$ in the obvious way. A unipotent automorphism of a mixed extension is an automorphism respecting the filtration and acting as the identity on the associated graded (i.e. an endomorphism of mixed extensions). of $V_{0}\oplus\ldots\oplus V_{n}$ as a $\mathbb{Q}_{p}$-vector space, and we have a bijection
@@ 2.2. Biextension structure
If $E_{1}$ and $E_{2}$ are extensions of $V_{0}$ by $V_{1}$ and $V_{1}$ by $V_{2}$ respectively, we will say that a mixed extension $M$ with graded pieces $V_{0},V_{1},V_{2}$ such that $\pi_{1}(M)\simeq E_{1}$ and $\pi_{2}(M)\simeq E_{2}$ is a mixed extension of $E_{1}$ and $E_{2}$ (where $\pi_{i}$ are the projection maps).
If we fix $E_{1}$, then the set of all mixed extensions in $M(V_{0},V_{1},V_{2})$ with $\pi_{1}(M)=[E_{1}]$ is in bijection with the group $\mathrm{Ext}^{1}(E_{1},V_{2})$. Similarly, if we fix $E_{2}$ then the set of all mixed extensions with $\pi_{2}(M)=[E_{2}]$ is in bijection with the group $\mathrm{Ext}^{1}(V_{0},E_{2})$. If $M_{1}$ and $M_{2}$ are two mixed extensions with $\pi_{i}(M_{1})=\pi_{i}(M_{2})$ ($i=1$ or $2$) then we will denote this induced addition by $M_{1}+_{i}M_{2}$.
@@
Note that $M(G;V_{0},V_{1},V_{2})$ does not in general satisfy the axioms of a bi-extension of abelian groups in the sense of [Mum69] (or [SGA72]), since we do not assume that the map
@@
is surjective. Indeed the possible failure of surjectivity is the main topic of this paper.

### v2

@@ Abstract.
The Chabauty–Coleman–Kim method in depth two describes the rational points on a curve in terms of a generalisation of Nekovář’s $p$-adic height pairing which replaces $\mathbb{G}_{m}$ with a higher Chow group. It is unclear both what the domain of definition of this pairing is, and how to compute it. This paper explores the relevance of the Beilinson–Bloch conjectures to this problem. In particular, it is shown that if $X$ is a smooth projective curve and the Albanese kernel of $X\times X$ is torsion, then there is an algorithm to compute the generalised height pairing on a pair of rational points on the Jacobian. This leads to the consideration of certain ‘motivic refinements’ of the nonabelian cohomology varieties which arise in nonabelian Chabauty.
@@ 1. Introduction
In Kim’s nonabelian generalisation of the Chabauty–Coleman method, one is instead interested in iterated extensions (henceforth referred to as mixed extensions) of geometric Galois representations. The graded pieces of these iterated extensions are quotients of tensor powers of the Tate module of the Albanese of the original variety $X$. The utility of working with this space of iterated extensions is that, while the space of geometric extensions of $\mathbb{Q}_{p}$ by $H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(1))$ may be large, conjectures on the dimensions of global Galois cohomology groups predict the space of geometric extensions of $\mathbb{Q}_{p}$ by higher tensor powers are ‘small’, and in particular most local iterated extensions do not come from global iterated extensions, giving an obstruction to a $p$-adic point coming from a rational point.
More precisely, let $b\in X(\mathbb{Q})$ and let $U_{n}$ be the maximal $n$-unipotent quotient of the $\mathbb{Q}_{p}$-unipotent completion of the étale fundamental group of $X_{\overline{\mathbb{Q}}}$ with basepoint $b$. Then one defines $X(\mathbb{Q}_{p})_{n}\subset X(\mathbb{Q}_{p})$ to be the set of points extending to an adelic point $(x_{v})\in X(\mathbb{A})$ whose class in $\prod_{v}H^{1}(\mathbb{Q}_{v},U_{n})$ lies in the image of $H^{1}(\mathbb{Q},U_{n})$. More generally, for any Galois stable quotient $U$ of $U_{n}$, one can similarly define a subset $X(\mathbb{Q}_{p})_{U}$ of $X(\mathbb{Q}_{p})$.
@@ 1.1. Applications to quadratic Chabauty
In full-fat quadratic Chabauty, one assigns to each pair of rational points on a smooth projective variety $X$ a Galois representation which is filtered with graded pieces $\mathbb{Q}_{p},V:=H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(1))$ and $\mathrm{Coker}(H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p})^{\otimes 2}\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}H^{2}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}))^{*}$. We write $M_{g}(\mathbb{Q};\mathbb{Q}_{p},V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$ for the set of such iterated extensions, with suitable local conditions (see §2.1 for details). This is a nonabelian cohomology variety which is slightly larger than the Selmer scheme considered by Kim. However, working with this larger space has benefits: it behaves more like the Selmer scheme of a elliptic curve minus the origin with tangential basepoint studied by Kim in [Kim10]. The reciprocity law giving a condition for a collection of local mixed extensions to come from a global mixed extension is then the generalised height pairing introduced in [BD21]. This has been used in [BD21] to determine rational points on bihyperelliptic curves of the form $y^{2}=x^{6}+ax^{4}+ax^{2}+1$, and in [Dog23] to determine the rational points on the hyperelliptic curve $y^{2}-y=x^{5}-x$.
@@
To explain why this is a difficult problem, consider the cases of classical Chabauty–Coleman and skimmed quadratic Chabauty. In the case of classical Chabauty, one determines the Chabauty–Coleman locus $X(\mathbb{Q}_{p})_{1}$ by finding a set of zero-cycles on $X$ generating a finite index subgroup of $J(\mathbb{Q})$. In skimmed quadratic Chabauty, one needs to determine the values of the $p$-adic height pairing. By analogy with the abelian case, one might hope that one can do this by evaluating the $p$-adic height function on points over number fields. However it is far from obvious that this works: for example in the case of integral points on a rank == genus hyperelliptic curve, one can determine the $p$-adic height pairing from $K$-points on $X$ (for $K/\mathbb{Q}$ a Galois extension, possibly infinite) if and only if $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$ is contained in the subgroup of $(\mathrm{Sym}^{2}J(K)\otimes\mathbb{Q})^{\operatorname{Gal}(K|\mathbb{Q})}$ generated by $(x-\infty)^{2}$, for $x\in X(K)$ and $\infty$ a Weierstrass point.
In practice one instead computes the $p$-adic height pairing using a more general construction – e.g. using algorithms for computing local height pairings between divisors $D_{1}$ and $D_{2}$ with disjoint support. From the perspective of mixed extensions, this can be interpreted as constructing all mixed extensions with graded pieces $\mathbb{Q}_{p},V,\mathbb{Q}_{p}(1)$ coming from geometry via subquotients of Galois representations of the form $H^{1}_{\mathrm{et}}(X_{\overline{\mathbb{Q}}}-|D_{1}|;|D_{2}|)(1)$ (see [Nek93], [Sch94]).
@@ 1.2. The Albanese kernel
Let $h$ denote the generalised height on mixed extensions with graded pieces $\mathbb{Q}_{p},V$ and $W:=\wedge^{2}V/\mathrm{NS}(J_{\overline{\mathbb{Q}}})^{*}\otimes\mathbb{Q}_{p}(1)$ (see section 2). This may be viewed either as a function on $X(\mathbb{Q})$, or as a function on $H^{1}_{f}(\mathbb{Q},V)\times H^{1}_{f}(\mathbb{Q},V^{*}\otimes W)$, or as a function on $J(\mathbb{Q})\times\mathrm{CH}^{2}(X^{3})_{0}$. As explained below, the relation between these is via the cycle class map
$\mathrm{Jac}(X)(K)\times\mathrm{CH}^{2}(X^{3})_{0}\to H^{1}(K,V)\times\to H^{1}(K,V^{*}\otimes W)$
@@
One of the original visions of the Chabauty–Coleman–Kim method (or “recurring fantasies”, in the language of the Beilinson–Bloch conjectures) was that the Bloch–Kato conjectures should imply the existence of an algorithm for determination of the rational points on curves [Kim12]. As is pointed out in loc. cit., by analogy with the Birch–Swinnerton-Dyer conjecture one would actually hope for an algorithm whose termination is conditional on these conjectures, meaning that one nevers needs to verify the Beilinson–Bloch conjectures for a variety in order to run the algorithm, but if the conjectures are false then the process is not guaranteed to terminate.
The present paper can be seen as an elaboration on one special case of that idea. [^2] With the caveat that Chabauty methods are essentially never known to give algorithms for determining rational points, because of well documented problems with possible nonrational points in $X(\mathbb{Q}_{p})_{n}$ or multiplicities of zeroes of the relevant $p$-adic power series (see [BS10] for a comprehensive description). Namely, we show that the problem of determining equations for $X(\mathbb{Q}_{p})_{2}$ essentially reduces to verifying (or more accurately, exhibiting) that certain elements of the Albanese kernel of $X\times X$ are torsion. To explain this,
@@ 1.3. Motivic avatars
(1) $1\to\mathbb{G}_{m}\to P^{\times}\to J\times J\to 1.$
@@
One can extend the Poincaré torsor to describe semi-skimmed quadratic Chabauty. One can think of this as involving working with $T$-torsors $\widetilde{P}^{\times}$ on $J$, or other abelian varieties, where $T$ is a nonsplit torus, or alternatively as simply working with the Poincaré torsor over number fields. However, beyond semi-skimmed quadratic Chabauty one would not expect any kind of motivic avatar of the Selmer scheme. A precise statement in this regard is the recent paper of Rogov which proves that when $n\geq 3$, Hain’s depth $n$ higher Albanese manifold (for a hyperbolic curve $X/\mathbb{C}$) is not algebraizable [Rog25]. An archetypal example of the kind of non-algebraic structure that can arise is that of ‘mixed Tate quadratic Chabauty’ – i.e. the study of $X(\mathbb{Z}_{p})_{2}$ when $X$ is $\mathbb{P}^{1}-\{0,1,\infty\}$. Here the relevant bilinear form comes from the $p$-adic dilogarithm function [DCW15]. Then there is a precise analogue of the exact sequence (1) defined in terms of the abelian category of mixed Tate motives. However, from a computational perspective, arguably a more useful analogue is the related (but slightly different) Bloch–Suslin complex.333perhaps a strict analogue would be a short exact of pointed sets constructed from the Bloch–Suslin complex, see 5.2. Instead of being short exact, this sequence has four terms, with the Milnor $K$-group $K^{M}_{2}(\mathbb{Q})$ providing an obstruction analogous to the obstruction coming from the Albanese kernel. In summary we have the following table of (sometimes imprecise) analogies.
@@
full-fat quadratic Chabauty $\left(\begin{array}[]{ccc}1&0&0\\ &\rho_{V}&0\\ &*&\rho_{V}^{\otimes 2}\end{array}\right)$ $\mathcal{M}(X^{2})$ $T(X^{2})$
@@
To elaborate briefly on the second point: Dan–Cohen, Wewers, Corwin [DCW15] [DCC18] and others have used the Tannakian category of mixed Tate motives to great effect in the study of Chabauty–Kim for the $S$-unit equation. Whilst the existence of such a category greatly clarifies structural properties, as far as the author is aware one still needs a geometric realisation of these mixed Tate motives in order to compute their $p$-adic periods. However, perhaps using an explicit description of the Galois group of mixed Tate motives in terms of algebraic cycles would give a practical approach to this computational problem [Blo91].
@@ Acknowledgements
It is a pleasure to be able to thank Minhyong Kim for the opportunity to spend so much time in the mathematical world he has created, and for his generosity and optimism. This paper benefited from many discussions with Carl Wang-Erickson about mixed extensions, and with Jonathan Love about the Albanese kernel.
@@ 2.1. Mixed extensions
Given objects $V_{0},\ldots,V_{n}$ in an abelian category $\mathcal{C}$, we can define a mixed extension with graded pieces $V_{0},\ldots,V_{n}$ to be an object $W$ of $\mathcal{C}$ together with a filtration by subobjects
$W=W_{0}\hookleftarrow W_{1}\hookleftarrow\ldots\hookleftarrow W_{n+1}=0$
together with isomorphisms $\psi_{i}:V_{i}\simeq W_{i}/W_{i+1}$. We define a morphism of mixed extensions $(W,(W_{i}),(\psi_{i}))\to(W^{\prime},(W^{\prime}_{i}),(\psi^{\prime}_{i}))$ to be a morphism $W\to W^{\prime}$ in $\mathcal{C}$, respecting the filtrations and commuting with the $\psi_{i}$ and $\psi_{i}^{\prime}$. If $\mathcal{C}$ is the category of continuous finite dimensional $\mathbb{Q}_{p}$-representations of a group $G$, then we can define a group $U(V_{0},\ldots,V_{n})$ of unipotent automorphisms444here we view $\oplus_{i=0}^{n}V_{i}$ as a mixed extension of the $V_{i}$ in the obvious way. A unipotent automorphism of a mixed extension is an automorphism respecting the filtration and acting as the identity on the associated graded (i.e. an endomorphism of mixed extensions). of $V_{0}\oplus\ldots\oplus V_{n}$ as a $\mathbb{Q}_{p}$-vector space, and we have a bijection
@@ 2.2. Biextension structure
If $E_{1}$ and $E_{2}$ are extensions of $V_{0}$ by $V_{1}$ and $V_{1}$ by $V_{2}$ respectively, we will say that a mixed extension $M$ with graded pieces $V_{0},V_{1},V_{2}$ such that $\pi_{1}(M)=[E_{1}]$ and $\pi_{2}(M)=[E_{2}]$ is a mixed extension of $E_{1}]$ and $E_{2}]$ (where $\pi_{i}$ are the obvious projection maps).
If we fix $E_{1}$, then the set of all mixed extensions in $M(V_{0},V_{1},V_{2})$ with $\pi_{1}(M)=[E_{1}]$ is in bijection with the group $\mathrm{Ext}^{1}(E_{1},V_{2})$. Similarly, if we fix $E_{2}$ then the set of all mixed extensions with $\pi_{2}(M)=[E_{2}]$ is in bijection with the group $\mathrm{Ext}^{1}(V_{0},E_{2})$. If $M_{1}$ and $M_{2}$ are two mixed extensions with $\pi_{i}(M_{1})=\pi_{i}(M_{2})$ ($i=1$ or $2$) then we will denote by $M_{1}+_{i}M_{2}$ the mixed extension obtained by Baer summation in this ext group.
@@
Note that $M(G;V_{0},V_{1},V_{2})$ does not in general satisfy the axioms of a bi-extension of abelian groups in the sense of [Mum69] or [SGA72], since we do not assume that the map
@@
is surjective. Indeed the possible failure of surjectivity is one of the main topics of this paper.